OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-8, 0) and a directrix at x = 8.

OpenStudy (anonymous):

@Hero

hero (hero):

Okay, go ahead and solve it using the formula I gave you previously. This time, the point for the directrix will be (8,y).

OpenStudy (anonymous):

working it meow

OpenStudy (anonymous):

I got \[y=x-8\sqrt{2}\]

OpenStudy (anonymous):

@Hero

hero (hero):

You should show the complete work you did on it. From what I observe, your final answer doesn't seem to have the right form.

OpenStudy (anonymous):

Im skipping the first step cause It takes so long to write \[\large x^2 + 64 +y^2 = x^2 - 64 +x^2 - y^2\]

OpenStudy (anonymous):

@Hero \[\large 64 = x^2 - y^2 - 64\] I feel like this is a formula of some sort...

hero (hero):

Skipping steps doesn't sound like such a good idea.

OpenStudy (anonymous):

I didn't skip any but the first where you plug in the values

OpenStudy (anonymous):

next i did this: \[\large y^2 = x^2 - 128\]

OpenStudy (anonymous):

Then took the square root of both sides

hero (hero):

That still doesn't look right.

hero (hero):

You must have made another mistake.

hero (hero):

BTW, \(y^2 + y^2 = 2y^2\)

OpenStudy (anonymous):

left that out by accident

OpenStudy (anonymous):

\[\large 64 = -2y^2 +x^2 -64\]

OpenStudy (anonymous):

Does that look right?

hero (hero):

I wrote out the steps by hand myself. There's a reason why I wanted you to post each step you did just like last time. If you skip steps, you end up confusing yourself. Also, if you never got to the point where you end up using difference of squares, then you may have made yet another mistake.

hero (hero):

I recommend starting over but posting each step this time

OpenStudy (anonymous):

Alright one sec

OpenStudy (anonymous):

\[\large (x+8)^2 +y^2 = (x-8)^2 + (x-y)^2\]

OpenStudy (anonymous):

\[\large then: x^2 +64 +y^2 = x^2 -64 +x^2 - y^2\]

OpenStudy (anonymous):

Do you see any problems yet? @hero

hero (hero):

Yes, I do actually. Double check your work.

OpenStudy (anonymous):

\[\large x^2 + 64 +y^2 = x^2 +64 +x^2 +y^2\]

OpenStudy (anonymous):

Is that correct?

hero (hero):



OpenStudy (anonymous):

oh Sh*t

OpenStudy (anonymous):

Can't believe I didn't see that earlier

hero (hero):

What didn't you see earlier?

OpenStudy (anonymous):

in the last parentheses in the formula I put an x instead of a y

OpenStudy (anonymous):

And when I simplify that down it comes out to \[\large y=\sqrt{-32x}\]

OpenStudy (anonymous):

unless I did that extremely wrong which I felt like I did

hero (hero):

Yes, because you should have isolated x instead of y.

OpenStudy (anonymous):

Ok. So what you had was correct

hero (hero):

The parabola is horizontal instead of vertical.

OpenStudy (anonymous):

Oh yeah XD Derp

OpenStudy (anonymous):

I have some more if you don't mind :)

OpenStudy (anonymous):

Different type of question

hero (hero):

Actually, I'm about to log off soon. Others should be available to help.

OpenStudy (anonymous):

ok thanks @hero